Consider the following two projects A and B. Assume that the appropriate discount rate for each project is 20%.

Year

CF for Project A

CF for Project B

0

1

2

3

- $100

70

80

90

- $1,700

900

900

900

If projects A and B are independent projects, which project(s) should you accept based on your best capital budgeting criteria? Please explain your rationale.
Answer #1

In independent projects, the acceptance of one project does not lead to the rejection of another project. Hence, we can choose multiple projects at the same time.

We accept project:

1. NPV method: if NPV is positive

2. IRR: if IRR > Discount rate

NPV is given by:

Σ Rt/(1 i)ttvaries from 1 to n NPV where Rt Cash flow netted (Inflow - Outflow) during the period t i Discount rate t numbe

For Project A

NPV = [70 / (1 + 20%)^1] + [80 / (1 + 20%)^2] + [90 / (1 + 20%)^3] - Initial Investment

NPV = 58.33 + 55.56 + 52.08 - 100

NPV = $ 65.97

For Project B

NPV = [900 / (1 + 20%)^1] + [900 / (1 + 20%)^2] + [900 / (1 + 20%)^3] - Initial Investment

NPV = 750 + 625 + 520.83 - 1700

NPV = $ 195.83

Since, NPV of both projects is positive, we can accept both the projects on the basis of NPV method

----------------------------------

2. IRR method:

The discount rate at which NPV = 0

For Project A

NPV = 0

NPV = [70 / (1 + r%)^1] + [80 / (1 + r%)^2] + [90 / (1 + r%)^3] - Initial Investment = 0

[70 / (1 + r%)^1] + [80 / (1 + r%)^2] + [90 / (1 + r%)^3] = 100

Solving for r,

r = 57.26%

For Project B

NPV = 0

NPV = [900 / (1 + r%)^1] + [900 / (1 + r%)^2] + [900 / (1 + r%)^3] - Initial Investment = 0

NPV = [900 / (1 + r%)^1] + [900 / (1 + r%)^2] + [900 / (1 + r%)^3] = 1700

Solving for r,

r = 27.24%

Since, IRR of both projects > Discount Rate (20%), we can accept both projects